Factorise the following expressions.
p^2 – 10p + 25 ?????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

p^2 – 10p + 25
= p^2-2×5×p+5^2
= (p-5)^2
Using identity:(x-y)2= x^2-2xy+y^2

Thanks